HX is a weak acid (Ka = 10-5). It forms a salt NaX (0.1M) on reacting with caustic soda. The degree of hydrolysis of NaX is

Grade:12

2 Answers

Aman Bansal
13 years ago

dear student,

Degree of hydrolysis of salt of weak acid with strong base = sqrt(Kw/(cKa))

Here Kw = 10^-14 Ka = 10^-5 c = 0.1

So dwgree of hydrolysis = sqrt(10^-8) = 10-4

BEST OF LUCK..!!!!

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

AMAN BANSAL                

Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
Degree of hydrolysis of salt of weak acid with strong base = (Kw/(cKa))
Here Kw = 10-14 Ka = 10-5 c = 0.1
So degree of hydrolysis = (10-8) = 10-4
 
Thanks and regards,
Kushagra